Question 16

ABC Company sells 38,436 chairs a year at an average price per chair of $23.28. The carrying cost per unit is $20.64. The company orders 133 chairs at a time and has a fixed order cost of $160.16 per order. The chairs are sold out before they are restocked. What are the total costs if the company orders the optimal quantity (Economic Order Quantity)?
Answer #1

EOQ = (2*Annual Demand*Cost per Order/Carrying cost per unit)^1/2

= (2*38436*160.16/20.64)^1/2

= 727.23 units i.e. 727 units

Number of orders = 38436/737 = 53 orders

Total cost = Purchase cost + Ordering cost + Carrying cost

= 38436*23.28 + 53*160.16 + (727/2)*20.64

= $910,781.2